OpenStudy (anonymous):

) Find the partial derivative indicated. Assume the variables are restricted to a domain on which the function is defined. d/dx (x^2e^(sqrt(2xy))

OpenStudy (anonymous):

@Brostep0s

OpenStudy (anonymous):

this is what i got x^2(e^(sqrt(2x))/(sqrt2x))+2xe^(sqrt(2xy))

OpenStudy (anonymous):

is that supposed to be d/dx times the rest?

OpenStudy (anonymous):

no its the derivative with respect to x i think

OpenStudy (michele_laino):

I think: \[\frac{ \delta }{ \delta x }x ^{2}e \sqrt{2xy}=\] \[=xe ^{\sqrt{2xy}}\left( 2+\frac{ x ^{2}y }{ \sqrt{2xy} } \right)\]

OpenStudy (michele_laino):

sorry \[x ^{2}e ^{\sqrt{2xy}}\]

OpenStudy (anonymous):

that didnt work out

OpenStudy (anonymous):

its assuming y is a constant i think, because it is the partial derivative

OpenStudy (michele_laino):

and: \[\frac{ \delta }{ \delta y }x ^{2}e ^{\sqrt{2xy}}=\frac{ x ^{3} }{ \sqrt{2xy} }e ^{\sqrt{2xy}}\]

OpenStudy (michele_laino):

I know that y is like a constant, please note that when you derive the function sqrt(2xy) you will get: \[\frac{ y }{ \sqrt{2xy} }\]

OpenStudy (anonymous):

d/dx with respect to x of sqrt(2xy)=2/(2(sqrt2))

OpenStudy (anonymous):

@myininaya i need your veteran expertise

OpenStudy (michele_laino):

@pmkat14 are you sure? and y where is, if you said it was a constant, then y will remain

OpenStudy (michele_laino):

oops where is y

OpenStudy (anonymous):

oh shoot your right

OpenStudy (anonymous):

so then it wouuld be x^2e^(y/sqrt(2xy))+(2xe^(sqrt(2xy))) ?

OpenStudy (michele_laino):

please, you say \[\frac{ \delta }{ \delta x }\]?

OpenStudy (anonymous):

\[2xe^\sqrt{2xy}+e ^{y/\sqrt{2xy}}x^2\]

OpenStudy (michele_laino):

please , I have made an error before so I'm very sorry. Now your first term is right! the second one it is no right!, because the exponent of e^(...) can not change

OpenStudy (michele_laino):

\[\frac{ y }{ \sqrt{2xy} }\] is the derivative of the exponent, which is a factor not an exponent!

OpenStudy (anonymous):

isnt the derivative of e^(any ecponent) ln(e)e^(original exponenet) where ln(e)=1

OpenStudy (michele_laino):

sorry, if you have to calculate the derivative of a function namely: \[f(x,y)=e ^{g(x,y)}\] then the derivative of f with respect to x is: \[e ^{g(x,y)}*\frac{ \delta g }{ \delta x }\] are you agree?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

then, please apply that above to calculate the second term

OpenStudy (michele_laino):

please note that in your case we have: \[g(x,y)=\sqrt{2xy}\] referring to the example above

OpenStudy (zarkon):

\[\partial\] \partial